OpenStudy (anonymous):

@UnkleRhaukus Hi I need help with some calculus questions! How do I find the derivative of s(t)=t^6 tan t - sqrt of t

OpenStudy (anonymous):

Chain Rule, Power Rule, Product Rule. These are the three you need to perform. Performing t\[\bf s'(t)=(t^6)'(\tan(t))+(\tan(t))'(t^6)-(\sqrt{t})'\]ese you will get:

OpenStudy (anonymous):

Can you help me step by step?

OpenStudy (unklerhaukus):

\[(x^n)'=nx^{n-1}\]

OpenStudy (unklerhaukus):

\[\sqrt x=x^{1/2}\]